OpenStudy (jaylelile):

5 times an integer is 20 more than 4 times the next consecutive even integer. what are the two integers?

OpenStudy (landon34):

first, lets write this into an equation.

OpenStudy (jaylelile):

how?

OpenStudy (landon34):

its easy. 5*x +20*4=n

OpenStudy (landon34):

Ill solve the equation for you.

OpenStudy (landon34):

multiply the 20 by 4: 20*4=80

OpenStudy (jaylelile):

ohhh thank you so much! so then do I multiply the 80 by 2 to get my next integer?

OpenStudy (landon34):

No, we reorder the terms. 80+5x = n

OpenStudy (landon34):

we add 80 to each side of the equation. 80 + -80 + 5x = -80 + n

OpenStudy (landon34):

then, we combine like terms: 80+80=0 0+5x=80+n

OpenStudy (landon34):

we divide each side by 5: x = -16 + 0.2n

OpenStudy (landon34):

finally, we simplify: x = 16 + 0.2n

OpenStudy (jaylelile):

ohhhhh okay thank you very much!

OpenStudy (landon34):

write these steps down, and study, you will get it if you study it.